OpenStudy (anonymous):

Which of the following is equivalent to log sub{9} 36? I got log36/log9 is this correct?

Parth (parthkohli):

Yep! :D

Parth (parthkohli):

\(\Rightarrow \log_xy =\Large {\log y \over \log x} \)

Parth (parthkohli):

So are certainly correct

OpenStudy (anonymous):

Yep, this is the change of base formula at work :-)

Parth (parthkohli):

you are*

OpenStudy (anonymous):

ha so do i get best answer lol kidding! ^.^

Parth (parthkohli):

Ohh 3 of em lol

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

@ParthKohli do you have some kind of formula clipboard open that you copy & paste from? That's so incredibly fast, no way you're typing all that syntax :-D

Parth (parthkohli):

I type it out @agentx5

OpenStudy (anonymous):

No way lol

Parth (parthkohli):

I have a 70 day experience lol

OpenStudy (anonymous):

:-P

OpenStudy (anonymous):

Thanks guys :)

Parth (parthkohli):

You're welcome, mustache girl :P

Parth (parthkohli):

*SHAKA*

OpenStudy (anonymous):

You wrote all this in a second?! \Rightarrow \log_xy =\Large {\log y \over \log x} I call shenanigans sir! :-3

Parth (parthkohli):

Yes. I did.

OpenStudy (anonymous):

^_^ You scary @ParthKohli

OpenStudy (anonymous):

haha

Parth (parthkohli):

Using equation editor takes 1000x than typing it out lol